16 lb of brand M cinnamon was made by combining 12 lb of Indonesian cinnamon which costs $19/lb with 4 lb of Thai cinnamon which costs $11/lb. Find the cost per lb of the mixture?
Answer by checkley75(3666)

19*12+11*4=16x
228+44=16x
16x=272
x=272/16
x=$17 is the cost of the mix.